OpenStudy (anonymous):

help? solve by compleating the square x²+10x=18 use the quadratic formula to solve x²+3=-4x

OpenStudy (anonymous):

\[x^2+10z=18\] \[(x+5)^2=18+25=43\] \[x+5=\pm\sqrt{43}\] \[x=-5\pm\sqrt{43}\]

OpenStudy (anonymous):

That isnt an option on my paper. are you sure?

OpenStudy (anonymous):

pretty sure, unless i made an arithmetic mistake \(18+25=43\) right?

OpenStudy (anonymous):

there shouldn't b a z in the beginning iif that changes anything...

Directrix (directrix):

@JaneDoe100 Have you done this problem yet? use the quadratic formula to solve x²+3=-4x

Directrix (directrix):

@JaneDoe100 The solutions of @satellite73 are correct. Are your options approximations? Or, in some other way?

OpenStudy (anonymous):

The options are 11.56 and 1.56 all i need 2 know is if one of them is negitive

OpenStudy (anonymous):

and we havent done that problem

OpenStudy (anonymous):

ah use a calculator then, to get \(-5+\sqrt{43}\) and \(-5-\sqrt{43}\)

OpenStudy (anonymous):

ok thanks! its -11.56 and +1.56 can you help me with the other one?

OpenStudy (anonymous):

@satellite73 ?